# AllSpark[Uncaptioned image]: A Multimodal Spatio-Temporal General Intelligence Model with Thirteen Modalities

### v2

@@ Abstract
RGB, infrared, multispectral, and other spatio-temporal modal data fundamentally represent different observational approaches for the same geographic object. Therefore, leveraging multimodal data is an inherent requirement for comprehending geographic objects. However, due to the high heterogeneity in structure and semantics among various spatio-temporal modal data, the joint interpretation of multimodal spatio-temporal data has long been an extremely challenging problem. The primary challenge resides in striking a trade-off between the cohesion and autonomy of diverse modalities. This trade-off becomes progressively nonlinear as the number of modalities expands. Inspired by the human cognitive system and linguistic philosophy, where perceptual signals from the five senses converge into language, we introduce the Language as Reference Framework (LaRF), a fundamental principle for constructing a multimodal unified model. Building upon this, we propose a multimodal spatio-temporal general artificial intelligence model called AllSpark. Our model integrates thirteen different modalities into a unified framework, including one-dimensional (text, code), two-dimensional (RGB, infrared, SAR, multispectral, hyperspectral, tables, graphs, trajectory, oblique photography), and three-dimensional (point clouds, videos) modalities. To achieve modal cohesion, AllSpark uniformly maps diverse modal features to the language modality by a modal bridge and a multimodal large language models. To maintain modality autonomy, AllSpark introduces modality-specific encoders to extract the tokens of various spatio-temporal modalities. Finally, observing a gap between the model‘s interpretability and downstream tasks, we designed modality-specific prompts and task heads to enhance the model’s generalization capability on specific downstream tasks. Experiments indicate that AllSpark, without expert knowledge of most spatio-temporal modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and trajectory compared to state-of-the-art models. Moreover, AllSpark shows excellent adaptability in various modalities, such as MSI, HSI, PointCloud, table, code, and graph. AllSpark demonstrates the potential and possibility of constructing general artificial intelligence with a large language model. LaRF and AllSpark contribute to the shift in the research paradigm in spatio-temporal intelligence, transitioning from a modality-specific and task-specific paradigm to a general paradigm.
@@ I Introduction
Figure 1: We propose ”Language as Reference Framework (LaRF)” as the fundamental principle for constructing multimodal models. Under the guidance of this principle, we introduced a multimodal general intelligence model, AllSpark, which integrates thirteen spatio-temporal modalities.
Benefiting from the increasingly diverse observational methods available for spatio-temporal scenes, geographic objects can be described using various spatio-temporal modalities, such as RGB, infrared, SAR, multispectral, graph, and spatio-temporal trajectory data. Each modality provides unique information about different aspects of the geographic object. Analogous to the human process of perceiving and understanding the world through multiple modalities, such as vision, hearing, and touch, joint interpretation of multimodal data is an inherent requirement for intelligent models to achieve cognition of geographic objects.
@@
For a long time, constrained by the high heterogeneity across various modalities mentioned above, researchers have often developed specific methods based on prior assumptions related to a particular modality or designed multimodal approaches based on a few low-heterogeneity modalities. For instance, in single-modal research, C. R. Qi et al. proposed PointNet[1] for the point cloud modality, emphasizing the invariance of point cloud data ordering and the significance of global and local features. For language modality, Vaswani et al. introduced the transformer[2], which focuses on the long-range dependencies within word sequences. For the graph modality, Kipf and Welling proposed the GCN[3] based on the adjacency relationships between nodes in a graph. In multimodal research, the fusion of optical and SAR imagery has been widely explored in both traditional and deep learning remote sensing[4]. Moreover, visual-language models have undergone rapid development in recent years[5, 6, 7, 8, 9]. The diverse prior assumptions associated with each modality have resulted in significant gaps between methods designed for different modalities, making it challenging to perceive and understand different modalities using a unified model.
@@
We believe that the key challenge in addressing this issue lies in striking a trade-off between the cohesion and autonomy of diverse modalities. Cohesion refers to the presence of mutually correlated shared information among modalities, while autonomy signifies the existence of unique information specific to each modality relative to others. If we simply project data from multiple modalities into a common representation space for fusion, emphasizing the cohesion between modalities, this approach may lead to the loss of modality-specific information, thereby diluting the essence of multimodal collaboration. In contrast, if we excessively stress the autonomy between modalities, it may hinder the establishment of connections among them, limiting the model’s ability to simultaneously perceive multiple modalities. Moreover, as the number of modalities increases, balancing cohesion and autonomy becomes progressively more challenging nonlinearly.
@@
We observe that in the process of comprehending the world, humans integrate information from multiple modalities, such as hearing, touch, smell, and vision. The concepts formed through the parsing of these modalities ultimately converge in language. Humans engage in associating, reasoning, and expressive behaviours through language. In other words, language precisely encodes human perception and understanding of the world, providing clear definitions and meanings to abstract concepts from each modality. Inspired by this, we propose the Language as Reference Framework (LaRF) as a fundamental principle for constructing multimodal models. As illustrated in Figure 1, the abstract concepts derived from each modality should align with language, allowing for joint interpretation in the unified representation space of language.
Building upon this, we propose a multimodal spatio-temporal general artificial intelligence model, AllSpark, that integrates thirteen different modalities into a unified framework, including one-dimensional (text, code), two-dimensional (RGB, infrared, SAR, multispectral, hyperspectral, tables, graphs, trajectory, oblique photography), and three-dimensional (point clouds, videos) modalities.
To achieve modal cohesion, AllSpark uniformly maps diverse modal features to the language modality. In addition, we designed modality-specific prompts to guide multimodal large language models in accurately perceiving multimodal data.
@@
To maintain the autonomy between modalities, AllSpark introduces specific modal encoders for each modality to extract independent tokens. Given the high heterogeneity among modality data and modality encoders, a significant dimensional gap exists between the tokens of each modality and the language modality. To address this issue, we introduce a modality bridge, a mechanism from Lynx[10], to accomplish dimensional mapping from each modality’s tokens to the language tokens.
@@
Finally, considering the existing gap between the interpretability of the multimodal large language model and the specific downstream tasks, we design task heads for each downstream task to enhance the model’s generalization capability. Given the powerful interpretability capabilities of the multimodal large language model, we adhere to a lightweight design principle in both the modal encoders and task heads.
@@
Experiments demonstrate that AllSpark achieves competitive accuracy without expert knowledge in most modalities and utilizing a unified structure in modalities such as RGB and spatio-temporal trajectories compared to state-of-the-art models. Specifically, in the RGB modality, the accuracy of AllSpark is only 0.84 lower than that of the SOTA model, and in the trajectory modality, the average displacement error (ADE) metric differs by only 0.07 compared to that of the SOTA model. Additionally, AllSpark exhibits excellent adaptability in various other modalities, including point clouds, multispectral, hyperspectral, tables, graphs, and code. Theoretically, our proposed model has the potential for seamless extension to an arbitrary number of modalities.
@@
  • We first propose a unified multimodal spatio-temporal general model, AllSpark, that successfully integrates thirteen spatio-temporal modalities into a single model.
  • Inspired by the human cognitive system and linguistic philosophy, we propose the language as reference framework (LaRF), which offers a novel solution to balance cohesion and autonomy among multiple modalities. This principle provides a fresh approach to multimodal unification.
  • We propose new approaches for designing multimodal models in the spatio-temporal domain. AllSpark demonstrates the potential and ability to construct a general model with large language models, thereby contributing to the shift in the research paradigm in spatio-temporal intelligence, transitioning from a modality-specific and task-specific paradigm to a general paradigm.
  • Experiments indicate that AllSpark, without expert knowledge of the majority of spatio-temporal modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and trajectory information compared to state-of-the-art models. Moreover, AllSpark shows excellent adaptability in various modalities, such as MSI, HSI, PointCloud, table, code, and graph. Theoretically, our proposed model has the potential for seamless extension to an arbitrary number of modalities.
@@ II Related Work
Leveraging multimodal data is an inherent requirement for achieving cognitive recognition of geospatial objects. An ideal multimodal model should possess the capability to integrate all the modalities for joint interpretation. Hence, a crucial trend in the research of intelligent methods in the spatio-temporal domain is the continual increase in the number of modalities available for joint interpretation. Initially, early researchers often constructed single-modal expert models based on prior assumptions about a specific modality, achieving remarkable success within each respective modality. In recent years, with a deeper understanding of single-modal interpretation methods, numerous researchers have attempted to integrate several low-heterogeneity modalities to construct multimodal interpretation approaches. However, as the number of modalities increases, the challenge of balancing cohesion and autonomy among the modalities becomes increasingly difficult. In the following, we recall the development of intelligent methods in the spatio-temporal domain from the perspective of the continually increasing number of modalities, and finally, we present the principles and approach of our proposed model, AllSpark.
@@ II-A Single-Modal Model
For one-dimensional modalities, we focus on code and text. Code and text are both fundamental language modalities, but due to the excellent characteristics of code, such as strict syntax, unambiguous nature, and ability to interact with machines, code is treated as separate modality. Z. Feng et al. pretrained a model, CodeBERT[11], which facilitates the mutual transfer of information between code and natural language modalities. For the text modality, landmark contributions include a transformer[2], BERT[12], and the GPT series [13, 14, 15, 16], which have inspired subsequent series of works.
We categorize RGB, multispectral, hyperspectral, infrared, SAR, table, graph, trajectory, and oblique photography as two-dimensional (2D) modalities.
Among the two-dimensional modalities, standard three-channel RGB images are among the most common. For the RGB modality, the ResNet[17] proposed by K. He et al., which is based on the importance of visual global and local information, and the visual transformer (ViT)[18] introduced by A. Dosovitskiy et al., which leverages a global attention mechanism, represent two landmark contributions.
@@
Beyond an increase in the number of bands, attention has been given to images captured in specific spectral bands. For instance, infrared images, widely applied in fields such as agriculture and oceanography, have gained significant attention. H. Chen et al. proposed a shallow convolutional neural network that integrates a decision tree to meet the demand for real-time intelligent interpretation of a large volume of infrared images in agricultural water resource monitoring[21]. In the case of SAR images formed by active microwave radar, S. Chen et al. introduced AConvNet[22], a widely used fully convolutional neural network for intelligent SAR image interpretation.
For oblique photography, 3D reconstruction aims to utilize single or multiple two-dimensional images from various viewpoints to reconstruct a three-dimensional model. Y. Yao et al. addressed this task by proposing MVSNet[23]. By utilizing techniques such as three-dimensional convolution, the model establishes correlations between images from different perspectives, thereby reconstructing a three-dimensional model of the target from the features extracted from two-dimensional images.
Beyond the visual domain, the table is one of the commonly used modalities for recording and expressing information. TabNet[24], proposed by S. Ö. Arik et al., employs a sequence attention mechanism to achieve feature selection in the Table modality, thereby enabling interpretable and more efficient learning.
@@
Finally, we turn our attention to three-dimensional modalities: video and point clouds.
The video modality is formed by stacking images over time. A model proposed by A. Karpathy et al. extends the connectivity of CNN models over time. Additionally, multiresolution techniques are utilized to accelerate the training process, effectively reducing the impact of spatio-temporal redundancy on model efficiency in the video modality[27].
@@ II-B Multimodal Model
For joint interpretation of two modalities, H. Li and X.-J. Wu recognized the enhancement possible through fusing RGB and infrared images, which convey information about colour, texture, and temperature. Given the widespread demand and application prospects in military, medical, and other fields, they proposed a dense fusion method for RGB and infrared images: DenseFuse[29]. This method, employing a dense module that connects the outputs of all convolutional layers, successfully achieves pixel-level fusion of the RGB and infrared modalities. A. Sadeghian et al. extended the social GANs[25] and introduced RGB images to enhance scene data in Sophie[30], achieving better results in trajectory prediction tasks. Recognizing the high complementarity between RGB and SAR images, L. H. Hughes et al. proposed a three-step deep neural network framework that utilizes a universal prediction of matching regions, generates heatmaps, and eliminates outliers to match RGB and SAR images[31]. X. Li et al. introduced the DTCDN[32], a model that employs a GAN network to migrate RGB and SAR images to the same feature space, facilitating target detection. J. Yang et al. proposed a dual-stream convolutional network that uses high-resolution multispectral images to enhance the spatial resolution of hyperspectral images[33].
@@
Notably, CLIP[5], proposed by A. Radford et al., associates the RGB modality with the text modality using contrastive learning. With pretraining guided by weak supervisory signals from text, CLIP has demonstrated outstanding capabilities in both visual single-modal tasks and visual-language multimodal tasks, inspiring a series of subsequent works[6, 7, 8, 9]. Y. Zhang et al. introduced a meta-transformer[34], leveraging the contrastive learning paradigm from CLIP to pretrain a universal backbone network under the visual-language modality. It exhibits multimodal generalization abilities across various modalities, such as point clouds, infrared, and hyperspectral data. J. Han et al. directly employed a multimodal large language model as a universal backbone network, proposing the One-LLM, which successfully unifies eight modalities, namely, images, audio, videos, and points[35]. The success of these approaches implies the unique role of language modalities in multimodal models.
@@
Therefore, building upon the aforementioned efforts, we systematically propose the fundamental principle of the LaRF. Guided by this principle, we balance cohesion and autonomy among diverse modalities and introduce a general intelligent model named AllSpark, which unifies thirteen spatio-temporal modalities and possesses the potential to extend to an arbitrary number of modalities.

### v3

# AllSpark[Uncaptioned image]: A Multimodal Spatio-Temporal General Intelligence Model with Ten Modalities via Language as a Reference Framework

@@ Abstract
RGB, multispectral, point and other spatio-temporal modal data fundamentally represent different observational approaches for the same geographic object. Therefore, leveraging multimodal data is an inherent requirement for comprehending geographic objects. However, due to the high heterogeneity in structure and semantics among various spatio-temporal modalities, the joint interpretation of multimodal spatio-temporal data has long been an extremely challenging problem. The primary challenge resides in striking a trade-off between the cohesion and autonomy of diverse modalities. This trade-off becomes progressively nonlinear as the number of modalities expands. Inspired by the human cognitive system and linguistic philosophy, where perceptual signals from the five senses converge into language, we introduce the Language as Reference Framework (LaRF), a fundamental principle for constructing a multimodal unified model. Building upon this, we propose AllSpark, a multimodal spatio-temporal general artificial intelligence model. Our model integrates ten different modalities into a unified framework, including one-dimensional (language, code, table), two-dimensional (RGB, SAR, multispectral, hyperspectral, graph, trajectory), and three-dimensional (point cloud) modalities. To achieve modal cohesion, AllSpark introduces a modal bridge and multimodal large language model (LLM) to map diverse modal features into the language feature space. To maintain modality autonomy, AllSpark uses modality-specific encoders to extract the tokens of various spatio-temporal modalities. Finally, observing a gap between the model’s interpretability and downstream tasks, we designed modality-specific prompts and task heads, enhancing the model’s generalization capability across specific tasks. Experiments indicate that the incorporation of language enables AllSpark to excel in few-shot classification tasks for RGB and point cloud modalities without additional training, surpassing baseline performance by up to 41.82%. Additionally, AllSpark, despite lacking expert knowledge in most spatio-temporal modalities and utilizing a unified structure, demonstrates strong adaptability across ten modalities. LaRF and AllSpark contribute to the shift in the research paradigm in spatio-temporal intelligence, transitioning from a modality-specific and task-specific paradigm to a general paradigm. The source code is available at https://github.com/GeoX-Lab/AllSpark.
@@ I Introduction
TABLE I: AllSpark integrates ten spatio-temporal modalities
Model 1D 2D 3D
Language Code Table RGB MSI HSI SAR Traj Graph Point
[…]
Figure 1: AllSpark demonstrates excellent adaptability across up to 10 heterogeneous modalities and shows outstanding few-shot learning capabilities in RGB and point cloud modalities.
Benefiting from the increasingly diverse observational methods available for spatio-temporal scenes, geographic objects can be described by various spatio-temporal modalities, such as RGB, SAR, multispectral, graph, point cloud and trajectory data [7, 8, 9]. Each modality provides unique information about different aspects of the geographic object. Analogous to the human process of perceiving and understanding the world through multiple modalities, such as vision, hearing, and touch, joint interpretation of multimodal data is an inherent requirement for intelligent models to achieve cognition of geographic objects.
@@
For a long time, constrained by the high heterogeneity across various modalities mentioned above, researchers have often developed specific methods based on prior assumptions related to a particular modality or designed multimodal approaches for a few low-heterogeneity modalities. For instance, in single-modal research, C. R. Qi et al. proposed PointNet[10] for the point cloud modality, emphasizing the invariance of point cloud data ordering and the significance of global and local features. For language modality, Vaswani et al. introduced the transformer[11], which focuses on the long-range dependencies within word sequences. For the graph modality, Kipf and Welling proposed the GCN[12, 13] based on the adjacency relationships between nodes in a graph. In multimodal research, the fusion of optical and SAR imagery has been widely explored in both traditional and deep learning remote sensing[14]. Moreover, visual-language models have undergone rapid development in recent years[15, 16, 17, 18, 19]. The diverse prior assumptions associated with each modality have resulted in significant gaps between methods designed for different modalities, making it challenging to perceive and understand different modalities using a unified model.
We believe that the key challenge in addressing this issue lies in striking a trade-off between the cohesion and autonomy of diverse modalities. In our paper, ”cohesion” refers to the presence of mutually correlated shared information among modalities. For instance, both RGB and SAR images may describe the contour of the same object. ”Autonomy,” on the other hand, refers to the existence of unique information specific to each modality relative to others. For example, the RGB modality can describe an object’s color and texture, while the SAR modality can capture the object’s scattering properties in relation to radar waves. Cohesion forms the foundation for the interrelation between modalities, while autonomy highlights the value of multimodal joint interpretation—gaining a complete understanding of an object by integrating multiple modalities.
@@
If we merely project data from different modalities into a shared representation space to emphasize inter-modality cohesion, this approach risks losing modality-specific information, ultimately undermining the unique contributions of each modality and weakening the core value of multimodal collaboration. In contrast, if we excessively stress the autonomy between modalities, it may hinder the establishment of connections among them, limiting the model’s ability to simultaneously perceive multiple modalities. Moreover, as the number of modalities increases, balancing cohesion and autonomy becomes progressively more challenging nonlinearly.
@@
We observe that in the process of comprehending the world, humans integrate information from multiple modalities, such as hearing, touch, smell, and vision. The concepts formed through the parsing of these modalities ultimately converge in language. Humans engage in associating, reasoning, and expressive behaviours through language. In other words, language precisely encodes human perception and understanding of the world, providing clear definitions and meanings to abstract concepts from each modality. Inspired by this, we propose the Language as Reference Framework (LaRF) as a fundamental principle for constructing multimodal models. It means that the abstract concepts derived from each modality should align with language, enabling joint interpretation in the unified representation space of language.
Building upon this, we propose a multimodal spatio-temporal general intelligence model [20], AllSpark, that integrates ten different modalities into a unified framework, including one-dimensional (language, code, table), two-dimensional (RGB, SAR, multispectral, hyperspectral, graph, trajectory), and three-dimensional (point cloud) modalities. As shown in Table I and Figure 1, previous work either overlooked some important spatio-temporal modalities, such as hyperspectral and trajectory, or focused solely on natural images without considering remote sensing imagery. AllSpark covers a broader range of spatio-temporal modalities, such as multispectral, hyperspectral, graph, trajectory, and more, while demonstrating excellent few-shot learning capabilities and modality adaptability.
@@
To achieve modal cohesion, AllSpark uniformly maps diverse modal features to the language feature space. To maintain the autonomy between modalities, AllSpark introduces specific modal encoders for each modality to extract independent tokens. Given the high heterogeneity among modality data and modality encoders, a significant dimensional gap exists between the tokens of each modality and the language modality. To address this issue, we introduce a modality bridge, a mechanism from perceiver[21], to accomplish dimensional mapping from each modality’s tokens to the language tokens [22].
@@
Finally, considering the existing gap between the interpretability of the multimodal large language model and the specific downstream tasks, we design task heads and modality-specific text prompts for each downstream task to enhance the model’s generalization capability. Given the powerful interpretability capabilities of the multimodal large language model, we adhere to a lightweight design principle in task heads.
@@
Experiments demonstrate that AllSpark, despite lacking expert knowledge in most spatio-temporal modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and spatio-temporal trajectories compared to state-of-the-art models. Specifically, in the RGB modality, the accuracy of AllSpark is only 0.84 lower than that of the SOTA model, and in the trajectory modality, the average displacement error (ADE) metric differs by only 0.07 compared to that of the SOTA model. Additionally, AllSpark exhibits excellent adaptability in various other modalities, including point cloud, multispectral, hyperspectral, table, graph, and code. Theoretically, our proposed model has the potential for seamless extension to an arbitrary number of modalities.
@@
  • We first propose a unified multimodal spatio-temporal general model, AllSpark, that successfully integrates ten spatio-temporal modalities into a single model.
  • Inspired by the human cognitive system and linguistic philosophy, we propose the language as reference framework (LaRF), which offers a novel solution to balance cohesion and autonomy among multiple modalities.
  • Experiments indicate that AllSpark demonstrates strong few-shot learning capabilities and supports ten modalities. Theoretically, our proposed model has the potential for seamless extension to an arbitrary number of modalities.
@@ II Related Work
Leveraging multimodal data is an inherent requirement for achieving cognitive recognition of geospatial objects. An ideal multimodal model should possess the capability to integrate all the modalities for joint interpretation. Hence, a crucial trend in the research of intelligent methods in the spatio-temporal domain is the continual increase in the number of modalities available for joint interpretation.
Initially, early researchers often constructed single-modal expert models based on prior assumptions about a specific modality, achieving remarkable success within each respective modality. In recent years, with a deeper understanding of single-modal interpretation methods, numerous researchers have attempted to integrate several low-heterogeneity modalities to construct multimodal interpretation approaches. However, as the number of modalities increases, the challenge of balancing cohesion and autonomy among the modalities becomes increasingly difficult.
In the following, we recall the development of intelligent methods in the spatio-temporal domain from the perspective of the continually increasing number of modalities, and finally, we present the principles and approach of our proposed model, AllSpark.
@@ II-A Single-Modal Model
For one-dimensional modalities, we focus on code, language and table. Given the excellent characteristics of code, such as strict syntax, unambiguous nature, and ability to interact with machines, code is treated as a separate modality.
Z. Feng et al. pretrained a model, CodeBERT[23], which facilitates the mutual transfer of information between code and natural language modalities. For the language modality, landmark contributions include a transformer[11], BERT[24], and the GPT series [25, 26, 27, 28], which have inspired subsequent series of works.
The table is one of the commonly used modalities for recording and expressing information. TabNet[29], proposed by S. Ö. Arik et al., employs a sequence attention mechanism to achieve feature selection in the Table modality, thereby enabling interpretable and more efficient learning.
We categorize RGB, multispectral, hyperspectral, SAR, graph, trajectory as two-dimensional (2D) modalities.
Among the two-dimensional modalities, standard three-channel RGB images are among the most common. For the RGB modality, the ResNet[30] proposed by K. He et al., which is based on the importance of visual global and local information, and the vision transformer (ViT)[31] introduced by A. Dosovitskiy et al., which leverages a global attention mechanism, represent two landmark contributions.
@@
In the case of SAR images formed by active microwave radar, S. Chen et al. introduced AConvNet[34], a widely used fully convolutional neural network for intelligent SAR image interpretation.
@@
Finally, we turn our attention to three-dimensional modalities: point cloud.
@@ II-B Multimodal Model
A. Sadeghian et al. extended the social GANs[35] and introduced RGB images to enhance scene data in Sophie[38], achieving better results in trajectory prediction tasks. Recognizing the high complementarity between RGB and SAR images, L. H. Hughes et al. proposed a three-step deep neural network framework that utilizes a universal prediction of matching regions, generates heatmaps, and eliminates outliers to match RGB and SAR images[39]. X. Li et al. introduced the DTCDN[40], a model that employs a GAN network to migrate RGB and SAR images to the same feature space, facilitating target detection. J. Yang et al. proposed a dual-stream convolutional network that uses high-resolution multispectral images to enhance the spatial resolution of hyperspectral images[41].
To achieve joint interpretation of multimodal data, traditional multimodal models typically design specific architecture based on the priors of certain modalities. For example, Hang et al. proposed Coupled CNNs [42], which consist of two CNN networks to extract spectral-spatial features from hyperspectral data and elevation information from LiDAR data. They ultimately use both feature-level and decision-level fusion methods to integrate the heterogeneous features of the two modalities. Similarly, Zhang et al. proposed SLA-Net [43], which designs specific network structures to extract and fuse spatial information and morphological characteristics from hyperspectral imagery. These methods are typically designed for specific purposes and modalities, making it difficult to extend them to more modalities.
Additionally, Gao et al. proposed DFINet [44], which extracts self-correlation and cross-correlation between multimodal data to deeply fuse hyperspectral and multispectral modality features. Likewise, Hong et al. introduced S2FL [45], which extracts Modality-Specific Subspaces for each modality and a Shared Subspace for all modalities, finally obtaining multimodal interpretation results through a unified projection. The issue with such methods is the lack of a unified alignment reference, limiting them to collaboration between a few modalities. As the number of modalities increases, the complexity of aligning multiple modalities will grow exponentially.
@@
Notably, CLIP[15], proposed by A. Radford et al., associates the RGB modality with the text modality using contrastive learning [46, 47]. With pretraining guided by weak supervisory signals from text, CLIP has demonstrated outstanding capabilities in both visual single-modal tasks and visual-language multimodal tasks, inspiring a series of subsequent works[16, 17, 18, 19]. Y. Zhang et al. introduced a meta-transformer[6], leveraging the contrastive learning paradigm from CLIP to pretrain a universal backbone network under the visual-language modality. It exhibits multimodal generalization abilities across various modalities, such as point cloud, infrared, and hyperspectral data. J. Han et al. directly employed a multimodal large language model as a universal backbone network, proposing the One-LLM, which successfully unifies eight modalities, namely, images, audio, videos, and points[5]. The success of these approaches implies the unique role of language modalities in multimodal models.
@@
Building upon the aforementioned efforts, we systematically propose the fundamental principle of the LaRF. Guided by this principle, we balance cohesion and autonomy among diverse modalities and introduce a general intelligent model named AllSpark, which unifies ten spatio-temporal modalities and possesses the potential to extend to an arbitrary number of modalities.